# MARRS: Masked Autoregressive Unit-based Reaction Synthesis

### v1

@@ 4 Experiments
Concatenate Fuse [34] $\underline{0.536^{\pm 0.0093}}$ $\underline{0.999^{\pm 0.00}}$ $10.635^{\pm 0.0562}$ $\underline{26.472^{\pm 0.1958}}$ $11.302^{\pm 0.7041}$ $0.743^{\pm 0.0002}$ $14.095^{\pm 0.1040}$ $\mathbf{24.405^{\pm 0.5096}}$
Cooperative Transformer [13] $0.589^{\pm 0.2348}$ $\mathbf{1.000^{\pm 0.00}}$ $\underline{10.60^{\pm 0.0470}}$ $26.291^{\pm 0.3186}$ $\underline{10.267^{\pm 0.2674}}$ $\underline{0.752^{\pm 0.0003}}$ $\mathbf{13.849^{\pm 0.1702}}$ $\underline{23.785^{\pm 0.2024}}$
@@
Table 3: Ablation studies on Action-Conditioned Fusion (ACF). “Concatenate Fuse” is a method proposed by ReGenNet [34] for a reactor to get information from an actor. “Cooperative Transformer” is a transformer structure designed by InterGen [13]. Our approach achieves SoTA on almost all metrics.
4.1 Experimental Setup
Datasets. Following [34], we evaluate our proposed framework on the NTU120-AS dataset using SMPL-X [20] models and actor-reactor annotations. NTU120-AS includes 8,118 human interaction sequences captured by 3 cameras across 26 action categories. As with [34], we select camera 1 and follow the cross-subject protocol, where half of the subjects are used for training and the remaining for testing.
Metrics. We employ the same evaluation metrics as [34], which are as follows: (1) Frechet Inception Distance (FID): measures the latent distribution distance between the generated dataset and the real dataset. (2) Accuracy of action recognition (Acc): assesses action-motion matching. (3) Diversity (Div): measures motion diversity in the generated motion dataset. (4) Multimodality (MultiMod): indicates diversity within the same active motion. We generate 1,000 samples 20 times with different random seeds and report the average along with the 95% confidence interval, following the methodology of previous works [27, 34].
@@ 4.1 Experimental Setup
4.2 Implementation Details.
For the first stage, the tokenizer consists of a 3-layer encoder-decoder architecture, featuring a hidden dimension of 256 and the batch size is set to 256. For the reaction generation, the transformer consists of 8 layers with a latent dim of 384, and the diffusion MLPs are composed of 3 layers, each with a hidden dimension of 1024. The AdamW optimizer [15] is utilized with $\beta_{1}=0.5$ and $\beta_{2}=0.99$ and batch size for this configuration is set to 128. For the training of all models, the learning rate attains 2e-4 following 1000 iterations, utilizing a linear warm-up schedule.
@@
Coordination Layer [41] $0.605^{\pm 0.0119}$ $\mathbf{1.000^{\pm 0.00}}$ $10.616^{\pm 0.0441}$ $26.297^{\pm 0.3194}$ $\underline{10.055^{\pm 0.4638}}$ $0.739^{\pm 0.0002}$ $14.178^{\pm 0.1677}$ $\underline{23.601^{\pm 0.2656}}$
@@
Table 4: Ablation studies on Adaptive Unit Modulation (AUM). “w.o. Unit Communication” indicates that the body and hands branches do not interact with each other. “Coordination Layer” is a method proposed by ParCo [41] for transferring information between different units. “Hands $\rightarrow$ Body” and “Body $\rightarrow$ Hands” represent a unidirectional transfer of information between units. Our approach achieves SoTA on almost all metrics.
4.3 Comparisons with State-of-the-arts.
4.3.1 Quantitative Results.
@@
Following [34], we mainly focus on the challenging and practical online, unconstrained setting. The overall results are shown in Tab. 1. In the train-conditioned setting, our method achieves state-of-the-art performance in terms of FID, action recognition accuracy, and multi-modality on the NTU120-AS dataset. This demonstrates that our model effectively learns the reaction patterns. As for the test-conditioned setting, our method achieves the best FID, action recognition accuracy, and multi-modality, especially FID, which is much better than other VAE-based and diffusion-based methods, which verifies that our method has better generalization ability.
4.3.2 Qualitative Comparisons.
@@
In Fig. 3, we provide a qualitative comparison between ReGenNet [34] and our proposed MARRS. It can be seen that the human reaction synthesized by our method has more reasonable relative positions and body movements. In particular, our method can generate more natural and plausible hand gestures due to modeling body and hands separately. We provide the supplemental demo named demo.mp4 in the supplementary materials.
4.4 Ablation Study
@@
In this section, we carry out extensive ablation experiments to investigate the effectiveness of the crucial components in MARRS, thus offering a more profound understanding of our method. Unless specified otherwise, we conduct all the ablations on the NTU120-AS dataset under the online, unconstrained setting.
4.4.1 Unit Division: Body and Hands
@@
In order to explore the necessity of dividing the whole-body into the body and hands units in the human reaction synthesis, we conduct the experiments on three methods: (i) keeping whole-body as a complete unit without unit division, (ii) dividing it into the upper-and-lower-body proposed by SCA [6], (iii) dividing it into six separate units proposed by ParCo [41]. Except for three different dividing ways, the training of the first stage VAE and the structure of the second stage model are exactly the same. The results are shown in Tab. 2. There is a significant decrease in FID, Acc, and MultiMod in the test-conditioned setting without unit-division. The performance of FID and MultiMod in the test-conditioned setting increase somewhat after applying the upper-and-lower-body division. When performing the 6-unit division, FID under train-conditioned and Acc under test-conditioned achieve the best results, however, FID under test-conditioned shows almost no improvement compared to w.o. unit division, and the number of division units increases making the training and inferring time slower. In contrast, our proposed approach achieves more balanced results under train-conditioned and test-conditioned, whereas Div and MultiMod under train-conditioned and FID under test-conditioned achieve the best performance.
4.4.2 Action-Conditioned Fusion
@@
To demonstrate the effectiveness of ACF, we additionally try two actor-reactor interactions, (i) concatenate fuse proposed by [34], and (ii) cooperative transformer proposed by [13]. The experimental results are shown in Tab. 3. Concatenate fuse only performs well on Multimod under test-conditioned, while the performance of all other metrics is unsatisfactory. After applying cooperative transformer, FID, Acc and Div under test-conditioned are improved to some extent. Compared with others, our proposed ACF obtains optimal performance in all four metrics under train-conditioned as well as FID and Acc under test-conditioned, proving that ACF can capture the motion patterns of the two persons well.
4.4.3 Adaptive Unit Modulation
@@
As noted in [41], communication between units is extremely important for generating stable whole-body motions. In this section, we delve into the effects of different communication ways on the model performance. We conduct the following sets of experiments: (i) no communication between different units, (ii) unit communication using the coordination layer proposed by ParCo [41], and (iii) only hands-to-body unidirectional information transfer using our proposed AUM, (iiii) only body-to-hands unidirectional information transfer using our proposed AUM. The experimental results are shown in Tab. 4. When there is no communication between body and hands units, the resulting metrics under train-conditioned and test-conditioned perform worse overall, yet there is no catastrophic decline [41] like in text-to-motion. And FID under both train-conditioned and test-conditioned was significantly improved after the use of the coordination layer, proving that the exchange of information between different units is still essential in human reaction synthesis. Next, we tried two unidirectional information transfer utilizing our proposed AUM. When only the hands unit passes information to the body unit, the model shows a significant degradation in some metrics, while when only the body unit passes information to the hands unit, the model performance improves overall, which demonstrates that using only the local information of the hands does not provide the body with sufficiently accurate positional information, and that instead, it is feasible to take advantage of the body’s information to optimize the details of the hands. Ultimately, the bidirectional communication of AUM achieves the best performance for all metrics under train-conditioned and FID, Acc, and Multimod under test-conditioned, demonstrating the effectiveness of our proposed AUM and the necessity of bidirectional unit communication.
@@
VQ-VAE $10.06^{\pm 0.32}$ $0.783^{\pm 0.0002}$
UD-VQ-VAE $\underline{5.76^{\pm 0.26}}$ $\underline{0.805^{\pm 0.0002}}$ $53.57^{\pm 6.35}$ $0.562^{\pm 0.0004}$
L2 Loss - - $\underline{15.95^{\pm 1.93}}$ $\mathbf{0.760^{\pm 0.0002}}$
@@
Table 5: Ablation studies on Autoregressive Reaction Diffusion. “VQ-VAE” is our re-implementation of the MMM [22] to human reaction tasks. “UD-VQ-VAE” is based on “VQ-VAE” by applying our proposed unit division (UD). “L2 Loss” means directly using L2 Loss to learn the corresponding continuous motion tokens instead of incorporating diffusion loss.
@@
cVAE [10] $74.73^{\pm 4.86}$ $0.760^{\pm 0.0002}$ $11.14^{\pm 0.04}$ $18.40^{\pm 0.26}$
AGRoL [4] $16.55^{\pm 1.41}$ $0.716^{\pm 0.0002}$ $13.84^{\pm 0.10}$ $21.73^{\pm 0.20}$
MDM [27] $7.49^{\pm 0.62}$ $\underline{0.775^{\pm 0.0003}}$ $\underline{13.67^{\pm 0.18}}$ $24.14^{\pm 0.29}$
MDM-GRU [27] $24.25^{\pm 1.39}$ $0.720^{\pm 0.0002}$ $\mathbf{13.43^{\pm 0.09}}$ $22.24^{\pm 0.29}$
ReGenNet [34] $6.19^{\pm 0.33}$ $0.772^{\pm 0.0003}$ $14.03^{\pm 0.09}$ $\mathbf{25.21^{\pm 0.34}}$
@@
Table 6: Results on the offline setting on NTU120-AS [34]. $\pm$ indicates 95% confidence interval, $\rightarrow$ means that closer to Real is better. Bold indicates the best and underline indicates second best.
@@
Table 7: Model scaling results and comparison of computational complexity. Our proposed MARRS can converge faster in the training process than ReGenNet [34] and enlarging the model size can enhance the overall generation performance.
4.4.4 Diffusion for Autoregressive Reaction Generation
Since autoregressive generation is usually accompanied by vector quantization (VQ), we first explore the performance of VQ-based methods on human reaction synthesis. As shown in Tab. 5, FID of the VQ-VAE-based method reaches 10.06 at the reconstruction stage, implying a failure at the reconstruction. Meanwhile, we apply the unit-division way in our proposed UD-VAE to VQ-VAE and name it UD-VQ-VAE, and the results of the reconstruction are significantly improved. However, FID and Acc in the generation phase only reached 53.57 and 0.562. In addition, we explore the use of UD-VAE in the first stage, while not incorporating the diffusion loss in the second stage, but directly using L2 Loss to learn the corresponding continuous motion tokens, which resulted in a drastic deterioration of FID, proving that the diffusion loss is necessary.
4.5 Extension to other settings
@@
To demonstrate the adaptability and generalization of our proposed method MARRS, we additionally provide the results of MARRS applied to the offline setting in Tab. 6. For the model, we simply adjust the attention mask in the transformer to fit the offline setting. As shown in Tab. 6, MARRS achieves optimal performance on FID and Acc and competitive results on Div and Multimod, highlighting the efficacy of our approach and its adaptability.
4.6 Model Scalability and Complexity
@@
To demonstrate the scaling ability of MARRS, we provide three versions of MARRS. These models range in size from around 30M, 60M, to 90M parameters. The details of the model performance, parameters, training time, and inference time are shown in Tab. 7. We note that enlarging the model size enhances the overall generation performance, particularly with respect to FID. Additionally, we provide a comparison of parameters, training time, and inference time for MARRS and ReGenNet [34]. It is evident that MARRS converges more rapidly during the training process than ReGenNet. Additionally, MARRS-Tiny enables faster inference while maintaining competitive performance.
5 Conclusion
In this paper, we introduce an innovative framework named MARRS, designed to generate synchronized and detailed reactions. Initially, we present the UD-VAE, which divides the whole body into distinct units: body and hands, allowing for independent encoding. Following this, we introduce ACF, a process that involves randomly masking reactive tokens and then extracting information of the body and hands from the remaining active tokens. To enable whole-body collaboration, we propose AUM, which uses information from either the body or hands to adaptively adjust the other unit. Finally, in our diffusion model, we utilize a compact MLP as a noise predictor for each specific unit and integrate diffusion loss to capture the probability distribution of each token. Both quantitative and qualitative evaluations indicate that our method outperforms existing methods.

### v2

@@ Introduction
Concatenate Fuse (Xu et al. 2024) $\underline{0.536^{\pm 0.0093}}$ $\underline{0.999^{\pm 0.00}}$ $10.635^{\pm 0.0562}$ $\underline{26.472^{\pm 0.1958}}$ $11.302^{\pm 0.7041}$ $0.743^{\pm 0.0002}$ $14.095^{\pm 0.1040}$ $\mathbf{24.405^{\pm 0.5096}}$
Cooperative Transformer (Liang et al. 2024) $0.589^{\pm 0.2348}$ $\mathbf{1.000^{\pm 0.00}}$ $\underline{10.60^{\pm 0.0470}}$ $26.291^{\pm 0.3186}$ $\underline{10.267^{\pm 0.2674}}$ $\underline{0.752^{\pm 0.0003}}$ $\mathbf{13.849^{\pm 0.1702}}$ $\underline{23.785^{\pm 0.2024}}$
@@
Table 3: Ablation studies on Action-Conditioned Fusion (ACF). “Concatenate Fuse” is a method proposed by ReGenNet (Xu et al. 2024) for a reactor to get information from an actor. “Cooperative Transformer” is a transformer structure designed by InterGen (Liang et al. 2024). Our approach achieves SoTA on almost all metrics.
Experimental Setup
Datasets. Following (Xu et al. 2024), we evaluate our framework on NTU120-AS and Chi3D-AS using SMPL-X (Pavlakos et al. 2019) models and actor-reactor annotations. NTU120-AS contains 8,118 interaction sequences from 26 actions, captured by 3 cameras. Following (Xu et al. 2024), we use camera 1 and the cross-subject protocol (half subjects for training, half for testing). Chi3D-AS results are in the supplementary materials.
Metrics. We employ the same evaluation metrics as (Xu et al. 2024), which are as follows: (1) Frechet Inception Distance (FID): measures the latent distribution distance between the generated dataset and the real dataset. (2) Accuracy of action recognition (Acc): assesses action-motion matching. (3) Diversity (Div): measures motion diversity in the generated motion dataset. (4) Multimodality (MultiMod): indicates diversity within the same active motion. We generate 1,000 samples 20 times with different random seeds and report the average along with the 95% confidence interval.
@@
Figure 3: Qualitative comparison with ReGenNet on human reaction synthesis. Blue for actors and Red for reactors.
@@
Coordination Layer (Zou et al. 2024) $0.605^{\pm 0.0119}$ $\mathbf{1.000^{\pm 0.00}}$ $10.616^{\pm 0.0441}$ $26.297^{\pm 0.3194}$ $\underline{10.055^{\pm 0.4638}}$ $0.739^{\pm 0.0002}$ $14.178^{\pm 0.1677}$ $\underline{23.601^{\pm 0.2656}}$
@@
Table 4: Ablation studies on Adaptive Unit Modulation (AUM). “w.o. Unit Communication” indicates that the body and hands branches do not interact with each other. “Coordination Layer” is a method proposed by ParCo (Zou et al. 2024) for transferring information between different units. “Hands $\rightarrow$ Body” and “Body $\rightarrow$ Hands” represent a unidirectional transfer of information between units. Our approach achieves SoTA on almost all metrics.
Comparisons with State-of-the-arts
Quantitative Results.
@@
Following (Xu et al. 2024), we mainly focus on the challenging and practical online, unconstrained setting. The overall results are shown in Tab. 1. In the train-conditioned setting, our method achieves state-of-the-art performance in terms of FID, action recognition accuracy, and multi-modality on the NTU120-AS dataset. This demonstrates that our model effectively learns the reaction patterns. As for the test-conditioned setting, our method achieves the best FID, action recognition accuracy, and multi-modality, especially FID, which is much better than other VAE-based and diffusion-based methods, which verifies that our method has better generalization ability. Offline setting results are in the supplementary materials.
Qualitative Comparisons.
@@
In Fig. 3, we provide a qualitative comparison between ReGenNet (Xu et al. 2024) and our proposed MARRS. It can be seen that the human reaction synthesized by our method has more reasonable relative positions and body movements. In particular, our method can generate more natural and plausible hand gestures due to modeling body and hands effectively. We provide more qualitative results in the supplementary materials.
Ablation Study
@@
In this section, we carry out extensive ablation experiments to investigate the effectiveness of the crucial components in MARRS, thus offering a more profound understanding of our method. Unless specified otherwise, we conduct all the ablations on the NTU120-AS dataset under the online, unconstrained setting. Model scalability and complexity details are in the supplementary materials.
Unit Division: Body and Hands.
@@
In order to explore the necessity of dividing the whole-body into the body and hands units in the human reaction synthesis, we conduct the experiments on three methods: (i) keeping whole-body as a complete unit without unit division, (ii) dividing it into the upper-and-lower-body proposed by SCA (Ghosh et al. 2021), (iii) dividing it into six separate units proposed by ParCo (Zou et al. 2024). Except for three different dividing ways, the training of the first stage VAE and the structure of the second stage model are exactly the same. The results are shown in Tab. 2. There is a significant decrease in FID, Acc, and MultiMod in the test-conditioned setting without unit-division. The performance of FID and MultiMod in the test-conditioned setting increase somewhat after applying the upper-and-lower-body division. When performing the 6-unit division, FID under train-conditioned and Acc under test-conditioned achieve the best results, however, FID under test-conditioned shows almost no improvement compared to w.o. unit division, and increasing the number of division units makes training and inference more cumbersome and slower. In contrast, our approach achieves more balanced results under train- and test-conditioned settings, while Div and MultiMod (train-conditioned) and FID (test-conditioned) achieve the best performance.
Action-Conditioned Fusion (ACF).
@@
To demonstrate the effectiveness of ACF, we additionally try two actor-reactor interactions, (i) concatenate fuse proposed by ReGenNet (Xu et al. 2024), and (ii) cooperative transformer proposed by InterGen (Liang et al. 2024). The experimental results are shown in Tab. 3. Concatenate fuse only performs well on Multimod under test-conditioned, while the performance of all other metrics is unsatisfactory. After applying cooperative transformer, FID, Acc and Div under test-conditioned are improved to some extent. Compared with others, our proposed ACF obtains optimal performance in all four metrics under train-conditioned as well as FID and Acc under test-conditioned, proving that ACF can capture the motion patterns of the two persons well.
Adaptive Unit Modulation (AUM).
@@
As noted in ParCo (Zou et al. 2024), communication between units is extremely important for generating stable whole-body motions. In this section, we delve into the effects of different communication ways on the model performance. We conduct the following sets of experiments: (i) no communication between different units, (ii) unit communication using the coordination layer proposed by ParCo (Zou et al. 2024), and (iii) only hands-to-body unidirectional information transfer using our proposed AUM, (iiii) only body-to-hands unidirectional information transfer using our proposed AUM. The experimental results are shown in Tab. 4. Without communication, metrics under train- and test-conditioned settings worsen but without catastrophic decline (Zou et al. 2024) seen in text-to-motion. And FID under both train-conditioned and test-conditioned was significantly improved after the use of the coordination layer, proving that the exchange of information between different units is still essential in human reaction synthesis. Next, we tried two unidirectional information transfer utilizing our proposed AUM. Using AUM unidirectionally, hands-to-body transfer degrades some metrics, while body-to-hands transfer improves overall performance. This demonstrates that using only the local information of the hands does not provide the body with sufficiently accurate positional information, and that, instead, it is feasible to leverage the body’s information to optimize the details of the hands. Ultimately, the bidirectional communication of AUM achieves the best performance for all metrics under train-conditioned and FID, Acc, and Multimod under test-conditioned, demonstrating the effectiveness of our proposed AUM and the necessity of bidirectional unit communication.
@@
VQ-VAE (AR) $10.06^{\pm 0.32}$ $0.783^{\pm 0.0002}$
UD-VQ-VAE (AR) $5.76^{\pm 0.26}$ $0.805^{\pm 0.0002}$ $53.57^{\pm 6.35}$ $0.562^{\pm 0.0004}$
ACTOR (VAE) $\underline{1.21^{\pm 0.23}}$ $\underline{0.823^{\pm 0.0002}}$ $20.16^{\pm 0.65}$ $0.730^{\pm 0.0004}$
AMDM (AR Diffusion) ${23.46}^{\pm{1.39}}$ ${0.718}^{\pm{.0002}}$
L2 Loss (VAE) $\mathbf{0.18^{\pm 0.10}}$ $\mathbf{0.866^{\pm 0.0001}}$ $\underline{15.95^{\pm 1.93}}$ $\mathbf{0.760^{\pm 0.0002}}$
@@
Table 5: Ablation studies on different framework. “VQ-VAE” is based on MMM (Pinyoanuntapong et al. 2024). “UD-VQ-VAE” is based on “VQ-VAE” by applying our proposed unit division (UD). “ACTOR” (Petrovich, Black, and Varol 2021) is a VAE-based generation method. “AMDM” (Shi et al. 2024) is an autoregressive (AR) diffusion method. “L2 Loss” means directly using L2 Loss instead of incorporating diffusion loss.
Diffusion for Autoregressive (AR) Reaction Generation.
Since AR generation is often accompanied by vector quantization (VQ), we first investigate the performance of VQ-based methods in human reaction synthesis. As shown in Tab. 5, VQ-VAE (AR) performs poorly (10.06 FID) in the first reconstruction phase, and the failed reconstruction means that the second generation phase is not feasible. Meanwhile, we apply the unit-division way in our proposed Unit Division to VQ-VAE and name it UD-VQ-VAE (AR). The results of the reconstruction are significantly improved, but it still performs poorly in the generation phase (53.57 FID). What is more, we apply ACTOR (VAE) (Petrovich, Black, and Varol 2021) and AMDM (AR Diffusion) (Shi et al. 2024) to human reaction, but neither FID nor accuracy achieves satisfactory performance. In addition, we explore not incorporating the diffusion loss in the second stage, but directly using L2 Loss to learn the corresponding continuous motion tokens, which resulted in a drastic deterioration of FID, proving that the diffusion loss is necessary.
Accuracy of Hand Poses and Global Translation
We used coordinate-based metrics (APE and AVE) to measure the accuracy of hand poses and global translation (root) with SOTA. As shown in Tab. 6, our method is more accurate in hand modeling, demonstrating the effectiveness of unit division and interaction. The error of global translation is also significantly smaller than that of ReGenNet, further highlighting the effectiveness of the MARRS framework.
[…]
ReGenNet ${0.420}^{\pm{.0001}}$ ${0.446}^{\pm{.0355}}$ ${0.327}^{\pm{.0001}}$ ${0.055}^{\pm{.0011}}$
MARRS(Ours) $\mathbf{{0.388}}^{\pm{.0001}}$ $\mathbf{{0.430}}^{\pm{.0357}}$ $\mathbf{{0.294}}^{\pm{.0001}}$ $\mathbf{{0.043}}^{\pm{.0010}}$
Table 6: Comparison on hands and global translation.
Conclusion
In this paper, we introduce an innovative framework named MARRS, designed to generate synchronized and fine reactions. Initially, we present the UD-VAE, which divides the whole body into distinct units: body and hands, allowing for independent encoding. Following this, we introduce ACF, a process that involves randomly masking reactive tokens and then extracting information of the body and hands from the remaining active tokens. To enable whole-body collaboration, we propose AUM, which uses information from either the body or hands to adaptively adjust the other unit. Finally, in our diffusion model, we utilize a compact MLP as a noise predictor for each specific unit and integrate diffusion loss to capture the probability distribution of each token. Both quantitative and qualitative evaluations indicate that our method outperforms existing methods.
Supplementary Material
Figure 4: Visualization of inference process. The generation of entire tokens is performed in an autoregressive manner. Compact diffusion model is very small, consisting of only a 3-layer MLP. Therefore, MARRS can achieve fast inference speed.
@@ Conclusion
cVAE (Kingma, Welling et al. 2013) $74.73^{\pm 4.86}$ $0.760^{\pm 0.0002}$ $11.14^{\pm 0.04}$ $18.40^{\pm 0.26}$
AGRoL (Du et al. 2023) $16.55^{\pm 1.41}$ $0.716^{\pm 0.0002}$ $13.84^{\pm 0.10}$ $21.73^{\pm 0.20}$
MDM (Tevet et al. 2022) $7.49^{\pm 0.62}$ $\underline{0.775^{\pm 0.0003}}$ $\underline{13.67^{\pm 0.18}}$ $24.14^{\pm 0.29}$
MDM-GRU (Tevet et al. 2022) $24.25^{\pm 1.39}$ $0.720^{\pm 0.0002}$ $\mathbf{13.43^{\pm 0.09}}$ $22.24^{\pm 0.29}$
ReGenNet (Xu et al. 2024) $6.19^{\pm 0.33}$ $0.772^{\pm 0.0003}$ $14.03^{\pm 0.09}$ $\mathbf{25.21^{\pm 0.34}}$
@@
Table 7: Results on the offline setting on NTU120-AS (Xu et al. 2024). $\pm$ indicates 95% confidence interval, $\rightarrow$ means that closer to Real is better. Bold indicates the best and underline indicates second best.
Method Train conditioned Test conditioned
[…]
Real $0.19^{\pm 0.01}$ $1.000^{\pm 0.000}$ $6.59^{\pm 0.21}$ $20.96^{\pm 0.46}$ $0.98^{\pm 0.22}$ $0.602^{\pm 0.007}$ $7.71^{\pm 0.29}$ $12.66^{\pm 0.35}$
MDM (Tevet et al. 2022) $12.64^{\pm 1.85}$ $\mathbf{1.000^{\pm 0.000}}$ $6.88^{\pm 0.24}$ $19.06^{\pm 0.29}$ $30.54^{\pm 3.49}$ $\underline{0.443^{\pm 0.007}}$ $6.51^{\pm 0.59}$ $10.23^{\pm 0.65}$
ReGenNet (Xu et al. 2024) $\underline{0.28^{\pm 0.01}}$ $\mathbf{1.000^{\pm 0.000}}$ $\underline{6.70^{\pm 0.13}}$ $\underline{20.75^{\pm 0.59}}$ $\underline{21.24^{\pm 3.42}}$ $0.442^{\pm 0.005}$ $\underline{6.96^{\pm 0.65}}$ $\underline{10.61^{\pm 0.90}}$
Ours $\mathbf{0.21^{\pm 0.01}}$ $\mathbf{1.000^{\pm 0.000}}$ $\mathbf{6.69^{\pm 0.14}}$ $\mathbf{20.90^{\pm 0.57}}$ $\mathbf{18.94^{\pm 2.21}}$ $\mathbf{0.480^{\pm 0.006}}$ $\mathbf{7.22^{\pm 0.53}}$ $\mathbf{11.13^{\pm 0.66}}$
Table 8: Comparison to state-of-the-arts on the online, unconstrained setting for human action-reaction synthesis on Chi3D-AS (Xu et al. 2024). $\pm$ indicates 95% confidence interval, $\rightarrow$ means that closer to Real is better. Bold indicates best result and underline indicates second best.
@@
Table 9: Model scaling results and comparison of computational complexity. Our proposed MARRS can converge faster in the training process than ReGenNet (Xu et al. 2024) and enlarging the model size can enhance the overall generation performance.
Implementation Details
For the first stage, the tokenizer consists of a 3-layer encoder-decoder architecture, featuring a hidden dimension of 256, and the batch size is set to 256. For the reaction generation, the transformer consists of 8 layers with a latent dimension of 384, and the diffusion MLPs are composed of 3 layers, each with a hidden dimension of 1024. The AdamW optimizer (Loshchilov and Hutter 2017) is utilized with $\beta_{1}=0.5$ and $\beta_{2}=0.99$ and the batch size for this configuration is set to 128. For the training of all models, the learning rate attains 2e-4 following 1000 iterations, utilizing a linear warm-up schedule.
Inference Process
Given the active motion sequence, we send it to the encoder of UD-VAE to get active tokens. We expect to generate the entire reactive tokens in $T$ iterations in an autoregressive manner, and the mask ratio in each iteration $t$ is determined by a decay scheduling function ${cos}(\frac{\pi t}{2T})$. First, all reactive tokens are masked. Active and reactive tokens are fed to ACF and AUM. Second, we take out the tokens that need to be unmasked from the predefined permuted sequences and send them into the compact diffusion (very small, 3-layer MLP) as conditions. Next, the tokens after sampling are regarded as the new reactive tokens, and the other tokens are masked again until the completion of the $T$ iterations. As for the diffusion procedure, we sample ${x}_{n}^{T}$ from a random Gaussian noise $\mathcal{N}(\mathbf{0},\mathbf{I})$. This initial sample is then progressively denoised in a step-by-step manner, transforming ${x}_{n}^{T}$ to ${x}_{n}^{0}$ through the iterative process defined by
$\begin{aligned} {x}_{n}^{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left({x}_{n}^{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}\left({x}_{n}^{t}|t,z_{n}\right)\right)+\sigma_{t}\epsilon,\end{aligned}$
where $\sigma_{t}$ represents the noise level at time step $t$, and $\epsilon$ is sampled from the Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$. Ultimately, the decoder in UD-VAE is used to decode all tokens and convert them back into reaction sequences. The overall inference flowchart is shown in Fig. 4.
More Qualitative Results
In Fig. 5, we provide a qualitative comparison between ReGenNet (Xu et al. 2024) and our proposed MARRS. It can be seen that the human reaction synthesized by our method has more reasonable relative positions and body movements. In particular, our method can generate more natural and plausible hand gestures due to modeling the body and hands separately. We have an additional supplemental video available named demo.mp4.
Offline Setting on NTU120-AS
@@
To demonstrate the adaptability and generalization of our proposed method MARRS, we additionally provide the offline setting results on NTU120-AS in Tab. 7. For the model, we simply adjust the attention mask in the transformer to fit the offline setting. As shown in Tab. 7, MARRS achieves optimal performance on FID and Acc and competitive results on Div and Multimod, highlighting the efficacy of our approach and its adaptability.
More Experiments on Chi3D-AS Dataset
To evaluate the generalizability of our approach, MARRS, we perform corresponding experiments on another dataset, Chi3D-AS (Xu et al. 2024). The open-source model of ReGenNet performs poorly on Chi3D-AS, and its results are different from those reported in the paper. However, the authors did not address this issue. Therefore, we re-train ReGenNet and the corresponding evaluation model using its open source code. We maintain the same experimental setup as described in the paper. The results are shown in Tab. 8. Our proposed MARRS outperforms both MDM (Tevet et al. 2022) and ReGenNet (Xu et al. 2024) across various metrics.
Model Scalability and Complexity
@@
To evaluate the scalability of MARRS, we present three versions of the model, with sizes of approximately 30M, 60M, and 90M parameters. Details regarding model performance, parameters, training time, and inference time are provided in Tab. 9. We observe that increasing the model size improves overall generation performance, particularly in terms of FID. Furthermore, we compare the parameters, training time, and inference time of MARRS with those of ReGenNet (Xu et al. 2024). The results indicate that MARRS converges more rapidly during training than ReGenNet. In addition, MARRS-Tiny achieves faster inference while maintaining competitive performance.
User Study
We conduct a user study to evaluate MARRS against ReGenNet (Xu et al. 2024). 20 users were asked to evaluate 32 samples. The results are shown in Fig. 6. Compared to the SOTA method, ReGenNet, about 76% of users believe that our motions are more natural, about 74% of users believe that our motions are smoother, and about 79% of users believe that our motions are more physically realistic.
Figure 6: User study. We use three subjective indicators, Naturalness, Smoothness, and Physical realism, to compare with the ReGenNet.
Limitation
Due to the lack of high-quality human reaction datasets, we validated the method only on NTU120-AS and Chi3D-AS. In addition, the generated motions exhibit slight foot sliding, which is not yet fully explored in this paper. Owing to the limited accuracy of the current dataset, the generation of certain fine-grained motions (such as finger contacts) is not sufficiently precise. Related researchers in the community are encouraged to explore more challenging aspects of the human reaction synthesis. We hope that MARRAS could provide some new insight for the community.